# Video Event Reasoning and Prediction by Fusing World Knowledge from LLMs with Vision Foundation Models

### v2

@@ 1 Introduction
The proliferation of video data has positioned it as a primary medium for information exchange and environmental perception, driving a significant evolution in computer vision research (fu2024survey; li2024multimodal). Historically, the field has achieved remarkable success in discriminative tasks—moving from foundational image recognition (dosovitskiy2021an) to complex action and gesture recognition in videos (abu-farha-2021-time). This progress is mirrored in specialized domains such as sign language translation, where hierarchical models have demonstrated powerful recognition capabilities , and in fine-grained sensing technologies like WiFi-based gesture and activity recognition, which have become increasingly robust and resistant to interference (gu2023attention; zhang2025wiopen). However, this paradigm of "recognition" primarily addresses the question of "what" is happening in a visual scene. The community is now shifting towards a more profound challenge: moving from mere perception to genuine cognition (zellers2019vcr). This involves enabling machines to reason about "why" an event is occurring and to predict "what" might happen next—a task that requires a level of understanding far beyond statistical pattern matching.
The primary obstacle to achieving this cognitive leap is the "knowledge gap." Current models, despite their architectural sophistication, often operate within a closed world, lacking the vast repository of commonsense, physical intuition, and social knowledge that humans effortlessly apply. For instance, while a model might accurately classify a video clip as "a person picking up an egg and flour," it typically fails to infer the underlying intent, such as "the person is about to bake a cake." This limitation becomes particularly salient when considering the demands of next-generation applications. Advanced human-machine systems (zhang2025wiopen), affective computing that aims to suppress label noise for reliability (zhang2025resup), and even emerging fields like psychological understanding via language models all necessitate a deeper, causal understanding of events. Similarly, while we can now capture nuanced physiological data through commodity devices (zhao2025wi), interpreting this data in the context of complex human activities requires advanced reasoning. The core challenge, therefore, is to imbue visual models with this external world knowledge.
@@
The remainder of this paper is organized as follows. Section 2 provides a comprehensive review of related literature. Section 3 presents the detailed architecture and components of our proposed methodology. Section 4 describes our experimental setup, datasets, and presents both quantitative and qualitative results. Finally, Section LABEL:sec:conclusion concludes the paper and discusses promising directions for future work.
@@ 2.3 The Rise of Vision-Language Models (VLMs)
The true catalyst for advanced visual reasoning was the effective fusion of vision and language. The development of CLIP (radford2021learning) demonstrated that a shared embedding space for images and text, learned via massive-scale contrastive pre-training, could enable remarkable zero-shot transfer capabilities. This breakthrough paved the way for a host of large-scale Vision-Language Models (VLMs). Early influential models like Flamingo (alayrac2022flamingo) introduced gated cross-attention layers to inject visual features into a pre-trained and frozen language model, showcasing impressive few-shot learning. This "frozen LLM" paradigm was further explored in works like (tsimpoukelli2021multimodal), which highlighted the potential of this frugal approach. Architectures like BLIP-2 (li2023blip) advanced this idea by introducing a lightweight "Q-Former" module to bridge the modality gap between a frozen image encoder and a frozen LLM, proving to be a highly effective and parameter-efficient strategy. The field continues to expand the scope of fusion, aiming to create omni-perception models like VALOR (chen2024valor) and LanguageBind (zhu2023languagebind), which align not just vision and text, but also audio, depth, and thermal data into a unified semantic space. This trend towards multimodal fusion is not limited to mainstream sensors; innovative research has shown the potential of fusing commodity WiFi signals with vision for tasks like emotion recognition (gu2023wife), illustrating a broader principle of synergistic sensing that our work builds upon.
@@ 2.6 Applications, Benchmarks, and Broader Context
Progress in this field is critically dependent on challenging and well-designed benchmarks. Datasets like CLEVRER (yi2020clevrer) specifically target causal and physical reasoning, while Test of Time (momeni2023test) focuses on evaluating temporal understanding. The Ego-Exo4D dataset (gandhi2023egoexo4d) pushes the frontier by providing synchronized first-person and third-person views of the same event, demanding a more holistic, cross-view understanding. While much research relies on traditional visual data, parallel advancements in alternative sensing modalities are creating new opportunities. Technologies using commodity WiFi (gu2023attention) and RFID (feng2025rf) are now capable of fine-grained activity and even keystroke detection (huang2024keystrokesniffer). These rich data streams, often generated in complex real-world settings like robotic vehicle perception (wang2023drive), require the same, if not more, sophisticated reasoning models to be interpreted meaningfully. This is especially true in healthcare, where vision-based assessment of Parkinson’s tremors (wu2024visionbased) or WiFi-based pulmonary function analysis (zhao2025wi) requires a deep understanding of subtle temporal patterns.
Finally, training these massive models on diverse, real-world data presents its own challenges, leading to research in areas like federated learning to handle decentralized data and heterogeneous networks, with frameworks like Finch enabling neural architecture search in such settings (liu2024finch). Our work is situated at the intersection of these advancements, aiming to leverage foundational models and advanced reasoning to create a system that not only understands video but can also anticipate its future, with broad implications for all these application domains.
@@ 4.1.1 Training and Evaluation Datasets
Finally, to rigorously assess the generalization capabilities of our framework, we reserve several challenging benchmarks exclusively for zero-shot evaluation. These include the Test of Time (ToT) benchmark (momeni2023test) to specifically diagnose temporal reasoning, the synthetic CLEVRER dataset (yi2020clevrer) to probe for emergent physical and causal intuition, and the VCR benchmark (zellers2019vcr) to evaluate commonsense reasoning on still frames derived from videos. None of these evaluation datasets were seen by the model during any phase of training.
@@ 4.2.3 Baselines for Comparison
4.3 Evaluation Metrics
Evaluating the nuanced tasks of reasoning and prediction requires a multi-faceted approach that goes beyond simple accuracy. We employ a combination of automatic metrics and human-centric evaluation.
4.3.1 Metrics for Closed-ended Tasks
For tasks formatted as multiple-choice questions (e.g., on NExT-QA and VCR), we report standard Accuracy, which measures the percentage of correctly answered questions.
4.3.2 Metrics for Open-ended Generation
For our primary tasks of generating free-form text for reasoning and prediction, we use a suite of metrics:
  • N-gram-based Metrics: We report standard corpus-based metrics including BLEU, ROUGE-L, METEOR, and CIDEr. These metrics measure the overlap of n-grams between the generated text and ground-truth references. While useful, they often fail to capture semantic correctness and logical coherence.
  • Embedding-based Metrics: To address the limitations of n-gram overlap, we use BERTScore, which computes the cosine similarity between the token embeddings of the generated and reference texts, providing a better measure of semantic similarity.
  • LLM-as-a-Judge: Recognizing that automated metrics are insufficient for evaluating the quality of reasoning, we adopt the "LLM-as-a-Judge" paradigm. We use GPT-4o as an impartial evaluator. A detailed prompt is crafted, asking the judge model to rate the generated response on a scale of 1 to 10 across three key axes: (1) Factual Grounding: Is the response consistent with the visual evidence in the video? (2) Logical Coherence: Is the reasoning sound and easy to follow? (3) Insightfulness: Does the response provide non-trivial insights or predictions? We report the average scores across a large, randomly sampled test set.
4.3.3 Human Evaluation
The ultimate arbiter of quality is human judgment. We conduct a human evaluation study on a subset of 200 randomly selected test instances. We present human annotators with the video and the outputs from our model and two strong baselines in a blind, randomized order. Annotators are asked to rank the outputs based on overall quality, considering correctness, coherence, and detail. We report the win/tie/loss percentages for our model against the baselines.
@@ 5.3 Ablation Studies
Example of Factual Hallucination Refer to caption
Prompt: "Describe the process shown in the video."
Our Model Output (with Hallucination): The mechanic is changing a tire. He loosens the lug nuts, jacks up the car, and removes the wheel. After mounting the new wheel and hand-tightening the nuts, he uses a torque wrench to ensure they are at the specified tightness. The final step is to lower the car.
Analysis: The model correctly identifies the overall process. However, it hallucinates the use of a torque wrench. While this is the correct professional procedure and reflects the model’s strong prior knowledge, the tool was not present or used in the actual video clip. This highlights the ongoing challenge of ensuring perfect factual grounding, where the LLM’s powerful priors can sometimes override direct visual evidence.
Figure 2: An example of a failure case (factual hallucination). Our model correctly describes the overall procedure but incorrectly introduces an object (a torque wrench) that was not present in the video, likely due to strong prior knowledge from its training data. This highlights the challenge of maintaining perfect visual grounding.

### v3

@@ 1 Introduction
The proliferation of video data has positioned it as a primary medium for information exchange and environmental perception, driving a significant evolution in computer vision research (fu2024survey; li2024multimodal). Historically, the field has achieved remarkable success in discriminative tasks—moving from foundational image recognition (dosovitskiy2021an) to complex action and gesture recognition in videos (abu-farha-2021-time). This progress is mirrored in specialized domains such as sign language translation, where hierarchical models have demonstrated powerful recognition capabilities , and in fine-grained sensing technologies like WiFi-based gesture and activity recognition, which have become increasingly robust and resistant to interference . However, this paradigm of "recognition" primarily addresses the question of "what" is happening in a visual scene. The community is now shifting towards a more profound challenge: moving from mere perception to genuine cognition (zellers2019vcr). This involves enabling machines to reason about "why" an event is occurring and to predict "what" might happen next—a task that requires a level of understanding far beyond statistical pattern matching.
The primary obstacle to achieving this cognitive leap is the "knowledge gap." Current models, despite their architectural sophistication, often operate within a closed world, lacking the vast repository of commonsense, physical intuition, and social knowledge that humans effortlessly apply. For instance, while a model might accurately classify a video clip as "a person picking up an egg and flour," it typically fails to infer the underlying intent, such as "the person is about to bake a cake." This limitation becomes particularly salient when considering the demands of next-generation applications. Advanced human-machine systems , affective computing that aims to suppress label noise for reliability, and even emerging fields like psychological understanding via language models all necessitate a deeper, causal understanding of events. Similarly, while we can now capture nuanced physiological data through commodity devices , interpreting this data in the context of complex human activities requires advanced reasoning. The core challenge, therefore, is to imbue visual models with this external world knowledge.
@@
The remainder of this paper is organized as follows. Section 2 provides a comprehensive review of related literature. Section 3 presents the detailed architecture and components of our proposed methodology. Section 4 describes our experimental setup, datasets, and presents both quantitative and qualitative results. Finally, Section concludes the paper and discusses promising directions for future work.
@@ 2.3 The Rise of Vision-Language Models (VLMs)
The true catalyst for advanced visual reasoning was the effective fusion of vision and language. The development of CLIP (radford2021learning) demonstrated that a shared embedding space for images and text, learned via massive-scale contrastive pre-training, could enable remarkable zero-shot transfer capabilities. This breakthrough paved the way for a host of large-scale Vision-Language Models (VLMs). Early influential models like Flamingo (alayrac2022flamingo) introduced gated cross-attention layers to inject visual features into a pre-trained and frozen language model, showcasing impressive few-shot learning. This "frozen LLM" paradigm was further explored in works like (tsimpoukelli2021multimodal), which highlighted the potential of this frugal approach. Architectures like BLIP-2 (li2023blip) advanced this idea by introducing a lightweight "Q-Former" module to bridge the modality gap between a frozen image encoder and a frozen LLM, proving to be a highly effective and parameter-efficient strategy. The field continues to expand the scope of fusion, aiming to create omni-perception models like VALOR (chen2024valor) and LanguageBind (zhu2023languagebind), which align not just vision and text, but also audio, depth, and thermal data into a unified semantic space. This trend towards multimodal fusion is not limited to mainstream sensors; innovative research has shown the potential of fusing commodity WiFi signals with vision for tasks like emotion recognition , illustrating a broader principle of synergistic sensing that our work builds upon.
@@ 2.6 Applications, Benchmarks, and Broader Context
Progress in this field is critically dependent on challenging and well-designed benchmarks. Datasets like CLEVRER (yi2020clevrer) specifically target causal and physical reasoning, while Test of Time (momeni2023test) focuses on evaluating temporal understanding. The Ego-Exo4D dataset (gandhi2023egoexo4d) pushes the frontier by providing synchronized first-person and third-person views of the same event, demanding a more holistic, cross-view understanding. While much research relies on traditional visual data, parallel advancements in alternative sensing modalities are creating new opportunities. Technologies using commodity WiFi and RFID are now capable of fine-grained activity and even keystroke detection . These rich data streams, often generated in complex real-world settings like robotic vehicle perception (wang2023drive), require the same, if not more, sophisticated reasoning models to be interpreted meaningfully. This is especially true in healthcare, where vision-based assessment of Parkinson’s tremors or WiFi-based pulmonary function analysis requires a deep understanding of subtle temporal patterns.
Finally, training these massive models on diverse, real-world data presents its own challenges, leading to research in areas like federated learning to handle decentralized data and heterogeneous networks, with frameworks like Finch enabling neural architecture search in such settings. Our work is situated at the intersection of these advancements, aiming to leverage foundational models and advanced reasoning to create a system that not only understands video but can also anticipate its future, with broad implications for all these application domains.